The sum of three numbers is 27 and their product is 585. Find the numbers.
Answer by Alan3354(69443)

The sum of three numbers is 27 and their product is 585. Find the numbers.
------------------
It's 3 unknowns and only 2 equations, not straightforward math.
Factor 585
585 = 3 x 3 x 5 x 13
13 has to be one of the numbers, because if it's multiplied by 3 or 5 the 27 is exceeded.
--------------
Now it's 2 numbers that add to 14 and the product is 45.
We can use algebra, but it's obvious they're 5 and 9.
------------
x + y = 14 --> x = 14-y
x*y = 45
------
sub for x
(14-y)*y = 45
y^2 - 14y + 45 = 0
(y-5)*(y-9) = 0
y = 5
y = 9
------------
5, 9 and 13